CAVITY RING DOWN SPECTROSCOPY OF MOLECULAR IONS IN THE 3 μm REGION

\underline{JOSEPH S. GUSS}, HARALD VERBRAAK and HAROLD LINNARTZ, Leiden Observatory, University of Leiden, 2300 RA Leiden, The Netherlands.

Ionic complexes may be thought of as intermediates whose potential energy surfaces will provide information on the reaction dynamics of molecule-ion reactions in space. Because ionic complexes possess significantly higher binding energies than their van der Waals analogues, they may survive the intense radiation conditions in space. Candidate complexes include (CO-CO)^{+}, H_{2}-HCO^{+}, H_{2}-HN_{2}^{+}, H_{2}O-H_{2}O^{+}, N_{2}-N_{2}^{+}, N_{2}-H^{+}-N_{2}. An experiment is described in which ionic complexes are created in a continuous plasma jet emanating from a slit-nozzle and characterized with a quadrupole mass spectrometer. Cavity ring down spectroscopy is used to map the potential energy surfaces and to determine molecular parameters. A cw optical parametric oscillator, operating in the 3 μm region is used as a light source. Presently, experiments are prepared to record fully resolved spectra of ionic complexes of astrophysical relevance, with a focus on species that may be considered reactive intermediates in ion-molecule reactions in space.